# Probing-RAG: Self-Probing to Guide Language Models in Selective Document Retrieval

### v1

@@ 1 Introduction
To optimize the RAG pipeline, recent research has introduced methods to adaptively adjust the number of retrievals. (Su et al., 2024; Mallen et al., 2022; Wang et al., 2023c). A notable example is Adaptive-RAG (Jeong et al., 2024), which classifies queries into three categories based on their complexity: no retrieval, single-step retrieval, or multi-step retrievals. However, as shown in Figure 1, external classifiers in Adaptive-RAG often fail to fully leverage the internal decision-making capabilities of the language model. This leads to unnecessary additional retrieval steps, resulting in knowledge conflicts between the model’s internal knowledge and externally retrieved information Xie et al. (2024); Xu et al. (2024). Other methods decide whether to retrieve based on token generation probabilities (Jiang et al., 2023) and linguistic feedback (Zhang et al., 2024; Ding et al., 2024) from the LLM. However, these approaches still rely solely on the final output and cannot fully capture the model’s internal reasoning processes.
@@ 2 Related Work
Adaptive retrieval methods (Jiang et al., 2023; Su et al., 2024; Asai et al., 2023) have emerged as an innovative approach that allows language models to actively decide when and what external information to retrieve based on the task type or the specific information contained within the query. These methods can be broadly categorized into three main approaches: external classifier based, LLM-based feedback, and confidence-based techniques.
@@
External classifier based methods focus on training external models to optimize retrieval decisions based on query characteristics. For instance, Adaptive-RAG (Jeong et al., 2024) incorporates an additional classifier that categorizes queries into three types: no retrieval, single-step retrieval, or multi-step retrieval. This classifier selects the optimal number of retrievals based on the complexity of each query. LLM-based feedback methods (Zhang et al., 2024; Ding et al., 2024) rely on the evaluation of response consistency to guide retrieval decisions. These methods use prompting or multilingual formulations to assess the quality of the model’s responses. If the LLM detects low consistency in its output, it triggers a retrieval to obtain relevant documents. Confidence-based approaches rely on the model’s token uncertainty to guide retrieval decisions. For example, FLARE (Jiang et al., 2023) initiates retrieval when any token in a generated sentence has a probability below a certain threshold, while DRAGIN (Su et al., 2024) considers both token uncertainty and the attention weights between successive tokens to determine when to retrieve additional information.
@@ 3.1 Prober
This normalized representation $T^{\prime\prime}_{l_{k}}$ is then fed into the prober:
@@ 3.2 Training Prober
Methods In-Domain Out-of-Domain Average
HotpotQA NQ TriviaQA MuSiQue 2Wiki
EM ACC EM ACC EM ACC EM ACC EM ACC EM ACC
No Retrieval 16.8 27.94 15.0 24.55 37.5 45.51 3.20 4.79 22.6 43.11 19.0 29.18
Single-step Approach 14.6 28.34 11.4 25.95 19.6 38.72 1.80 5.79 22.8 38.32 14.0 27.42
LLM-based 18.5 25.74 17.6 20.35 36.7 41.91 3.80 4.59 24.2 25.75 20.2 23.67
FLARE 13.2 20.96 9.00 21.76 13.8 30.94 1.20 5.00 21.6 27.74 11.8 21.28
DRAGIN 19.8 22.55 18.8 22.16 42.7 47.11 4.20 4.79 26.5 28.74 22.4 25.07
Adaptive-RAG 13.2 23.55 11.4 26.18 22.8 40.72 1.20 3.00 21.6 40.52 14.0 26.79
Probing-RAG(Ours) 21.8 39.32 21.6 34.93 41.7 52.29 4.80 8.78 24.2 43.51 22.8 35.77
@@ 3.3 Probing based Retrieval-Augmented Generation
After generating the initial $r$ and $\hat{a}$, the prober assesses whether retrieval is necessary. To do this, we extract hidden state representations $T^{\prime\prime}_{l_{k}}$ of the $r$ and $\hat{a}$ tokens from intermediate layers and feed them into the probers assigned to each layer to generate logit values. The decision to perform additional retrieval is based on the sum of the probers’ logit values. If the difference between the logit for retrieval necessity and the logit indicating no need for retrieval is higher than the threshold $\theta$, additional documents ${d_{k}}_{1}^{j}$ are retrieved based on the following conditions:
@@ 4.1 Experimental Setup
We use the following baselines to evaluate the performance: No Retrieval, which proceeds directly based on the query without retrieval. Single-step Approach, which conducts a single retrieval call before QA. The LLM-based approach determines whether retrieval is necessary based on the language model’s response to the question. FLARE (Jiang et al., 2023) is a multi-round retrieval-augmented method that triggers retrieval each time it encounters an uncertain token. When a sentence contains uncertain tokens, a query is generated to retrieve relevant passages. The sentence is then replaced with a newly generated one based on the query and the retrieved passage. DRAGIN (Su et al., 2024) determines when to retrieve based on token generation probabilities and performs query reformulation using attention weights. Adaptive RAG (Jeong et al., 2024) performs by fine-tuning a query complexity classifier. All methods use 4-shot examples, and we evaluate performance using Exact Match (EM) and Accuracy (ACC).
@@
We use BM25 (Robertson and Jones, 1976; Robertson et al., 2009), a term-based sparse retrieval model, for all frameworks to ensure a fair comparison. We use the same document corpus and datasets as Jeong et al. (2024). For Adaptive-RAG, we train the classification model using the t5-large (Chung et al., 2024) model, following the approach of Jeong et al. (2024). We use the Gemma-2b (Team et al., 2024) model as the QA model for all frameworks. We use the residual post (epoch=2) as the prober’s hidden state representation position for Probing-RAG training and evaluation.
@@ 4.2 Main Results
Table 2: Distribution of retrieval steps showing the proportion of instances with no, single-step, and multi-step retrieval.
@@ 4.3 Analysis
Table 3: Performance comparison of variations in training data size for Probing-RAG and different retrieval methods. Values shown are accuracy (%) for in-domain, out-of-domain, and total averages.
Figure 4: (Left) Probing accuracy measured for each model among the layer, (Right) Correlation between prober’s classification performance and QA performance.
@@
To analyze the impact of training data size on performance in our framework, we conduct experiments by training the prober with varying amounts of data. We measure the performance on open-domain QA by training the prober on reduced datasets consisting of a sample of 1,000 data, as well as 1/4 and 1/2 proportions of the total 26,060 data points present the result in Table 3. As expected, the overall results show that in most cases, performance improves as the amount of training data increases. However, we also observe that the model trained with only 1k data points is comparable to the prober trained on one-third of the entire dataset in most of the datasets. Furthermore, we find that the prober trained on just 1k data points outperforms all of the previous methods, indicating that that it is possible to effectively train the prober using a small dataset.
@@
Table 4: Comparison of accuracy between no retrieval and adaptive retrieval augmentation to evaluate model consistency.
@@ 4.4 Prober Accuracy
Figure 5: Kernel density estimate plot of logits, where orange indicates retrieval calls and blue indicates no retrieval needed. Marginal distributions are shown on the top and right. The results are projected onto the 10th and 12th residual post positions.

### v2

@@ 1 Introduction
To optimize the RAG pipeline, recent research has introduced methods to adaptively adjust the number of retrievals. (Su et al., 2024; Mallen et al., 2022; Wang et al., 2023c). A notable example is Adaptive-RAG (Jeong et al., 2024), which classifies queries into three categories based on their complexity: no retrieval, single-step retrieval, or multi-step retrievals. However, as shown in Figure 1, external classifiers in Adaptive-RAG often fail to fully leverage the internal decision-making capabilities of the language model. This leads to unnecessary additional retrieval steps, resulting in knowledge conflicts between the model’s internal knowledge and externally retrieved information Xie et al. (2024); Xu et al. (2024). Other methods decide whether to retrieve based on token generation probabilities (Jiang et al., 2023b) and linguistic feedback (Zhang et al., 2024; Ding et al., 2024) from the LLM. However, these approaches still rely solely on the final output and cannot fully capture the model’s internal reasoning processes.
@@ 2 Related Work
Adaptive retrieval methods (Jiang et al., 2023b; Su et al., 2024; Asai et al., 2023) have emerged as an innovative approach that allows language models to actively decide when and what external information to retrieve based on the task type or the specific information contained within the query. These methods can be broadly categorized into three main approaches: external classifier based, LLM-based feedback, and confidence-based techniques.
@@
External classifier based methods focus on training external models to optimize retrieval decisions based on query characteristics. For instance, Adaptive-RAG (Jeong et al., 2024) incorporates an additional classifier that categorizes queries into three types: no retrieval, single-step retrieval, or multi-step retrieval. This classifier selects the optimal number of retrievals based on the complexity of each query. LLM-based feedback methods (Zhang et al., 2024; Ding et al., 2024) rely on the evaluation of response consistency to guide retrieval decisions. These methods use prompting or multilingual formulations to assess the quality of the model’s responses. If the LLM detects low consistency in its output, it triggers a retrieval to obtain relevant documents. Confidence-based approaches rely on the model’s token uncertainty to guide retrieval decisions. For example, FLARE (Jiang et al., 2023b) initiates retrieval when any token in a generated sentence has a probability below a certain threshold, while DRAGIN (Su et al., 2024) considers both token uncertainty and the attention weights between successive tokens to determine when to retrieve additional information.
@@ 3.1 Prober
This normalized representation $T^{\prime}_{l_{k}}$ is then fed into the prober:
@@ 3.3 Probing based Retrieval-Augmented Generation
After generating the initial $r$ and $\hat{a}$, the prober assesses whether retrieval is necessary. To do this, we extract hidden state representations $T^{\prime}_{l_{k}}$ of the $r$ and $\hat{a}$ tokens from intermediate layers and feed them into the probers assigned to each layer to generate logit values. The decision to perform additional retrieval is based on the sum of the probers’ logit values. If the difference between the logit for retrieval necessity and the logit indicating no need for retrieval is higher than the threshold $\theta$, additional documents ${d_{k}}_{1}^{j}$ are retrieved based on the following conditions:
@@
Methods In-Domain Out-of-Domain Average
HotpotQA NQ TriviaQA MuSiQue 2Wiki
EM ACC EM ACC EM ACC EM ACC EM ACC EM ACC
Gemma-2b
   No Retrieval 16.8 28.0 15.0 24.6 37.4 45.4 3.2 4.8 22.6 43.0 19.0 29.2
   Single-step Approach 14.6 28.2 11.4 26.0 19.6 38.8 1.8 5.8 22.8 38.4 14.0 27.4
   LLM-based 18.6 25.8 17.6 20.4 36.8 41.8 3.8 4.6 24.2 25.8 20.2 23.7
   FLARE 13.2 21.0 9.0 21.8 13.8 31.0 1.2 5.0 21.6 27.8 11.8 21.3
   DRAGIN 19.8 22.6 18.8 22.2 42.8 47.0 4.2 4.8 26.4 28.8 22.4 25.1
   Adaptive-RAG 13.2 23.6 11.4 26.2 22.8 40.8 1.2 3.0 21.6 40.6 14.0 26.8
   Probing-RAG(Ours) 21.8 39.4 21.6 35.0 41.8 52.2 4.8 8.8 24.2 43.6 22.8 35.8
Mistral-7b
   No Retrieval 17.0 20.6 13.2 19.8 38.0 45.2 3.4 6.2 16.4 30.0 17.6 24.4
   Single-step Approach 18.6 34.2 16.8 35.0 34.6 51.0 5.4 9.0 21.6 32.6 19.4 32.4
   LLM-based 20.4 32.0 15.8 35.6 41.2 49.8 5.6 9.4 16.8 32.4 20.0 31.8
   FLARE 20.4 32.0 15.4 34.4 35.0 45.6 4.4 6.6 18.6 31.0 18.8 29.9
   DRAGIN 21.2 28.0 16.8 37.2 39.8 42.2 5.2 7.2 23.2 25.8 21.2 28.1
   Adaptive-RAG 19.0 26.0 17.2 37.4 40.8 50.2 4.0 5.8 22.6 31.6 20.7 30.2
   Probing-RAG(Ours) 22.4 38.6 20.8 39.4 43.2 52.2 5.8 9.8 23.0 33.4 23.0 34.7
@@ 4.1 Experimental Setup
We use the following baselines to evaluate the performance: No Retrieval, which proceeds directly based on the query without retrieval. Single-step Approach, which conducts a single retrieval call before QA. The LLM-based approach determines whether retrieval is necessary based on the language model’s response to the question. FLARE (Jiang et al., 2023b) is a multi-round retrieval-augmented method that triggers retrieval each time it encounters an uncertain token. When a sentence contains uncertain tokens, a query is generated to retrieve relevant passages. The sentence is then replaced with a newly generated one based on the query and the retrieved passage. DRAGIN (Su et al., 2024) determines when to retrieve based on token generation probabilities and performs query reformulation using attention weights. Adaptive RAG (Jeong et al., 2024) performs by fine-tuning a query complexity classifier. All methods use 4-shot examples, and we evaluate performance using Exact Match (EM) and Accuracy (ACC).
@@
We use BM25 (Robertson and Jones, 1976; Robertson et al., 2009), a term-based sparse retrieval model, for all frameworks to ensure a fair comparison. We use the same document corpus and datasets as Jeong et al. (2024). For Adaptive-RAG, we train the classification model using the t5-large (Chung et al., 2024) model, following the approach of Jeong et al. (2024). We use the Gemma-2b (Team et al., 2024) and Mistral-7b (Jiang et al., 2023a) model as the QA model for all frameworks. We use the residual post (epoch=2) as the prober’s hidden state representation position for Probing-RAG training and evaluation.
@@ 4.2 Main Results
Table 2: Distribution of retrieval steps using the Gemma-2b model, showing the proportion of instances with no, single-step, and multi-step retrieval.
@@ 4.3 Analysis
Table 3: Performance comparison of variations in training data size for Probing-RAG and different retrieval methods using the Gemma-2b model. Values shown are accuracy (%) for in-domain, out-of-domain, and total averages.
Figure 4: (Left) Probing accuracy measured for each model among the layer, (Right) Correlation between prober’s classification performance and QA performance, using the Gemma-2b model.
@@
To analyze the impact of training data size on performance in our framework, we conduct experiments by training the prober with varying amounts of data. We measure the performance on open-domain QA by training the prober on reduced datasets consisting of a sample of 1,000 data, as well as 1/4 and 1/2 proportions of the total 26,060 data points we present the result in Table 3. As expected, the overall results show that in most cases, performance improves as the amount of training data increases. However, we also observe that the model trained with only 1k data points is comparable to the prober trained on one-third of the entire dataset in most of the datasets. Furthermore, we find that the prober trained on just 1k data points outperforms all of the previous methods, indicating that that it is possible to effectively train the prober using a small dataset.
@@
Table 4: Comparison of accuracy between no retrieval and adaptive retrieval augmentation to evaluate model consistency using the Gemma-2b model.
@@ 4.4 Prober Accuracy
Correlation between Prober Accuracy and QA Performance
Figure 5: Kernel density estimate plot of logits using the Gemma-2b model, where orange indicates retrieval calls and blue indicates no retrieval needed. Marginal distributions are shown on the top and right. The results are projected onto the 10th and 12th residual post positions.